### v1

@@ 3.2.3 Training Strategies
User-LLM offers flexible training strategies to tailor its performance for diverse use cases. To thoroughly explore User-LLM’s capabilities, we primarily investigated four strategies targeting different model components:
@@
  • LoRA: Finetuning the LLM with LoRA (Low-Rank Adaptation) Hu et al. (2022), along with the user encoder and projection layers, offers parameter efficiency preventing catastrophic forgetting and maintaining the LLM’s core knowledge.
@@ 4.1 Datasets and Tasks
Dataset Train Examples Test Example Seq Len Item
MovieLens20M 13,821,405 81,970 50 Movie
Google Review 3,387,375 80,080 50 Place
Amazon Review 357,258 5,756 50 Product
Table 1: Dataset train / test split sizes, sequence lengths, and item features.
We evaluated our approach on three widely recognized datasets. Table 1 summarizes their statistics after processing. Notably, Amazon Review dataset is well known for its high sparsity and variability and contains a significantly smaller number of training and evaluation examples.
  • MovieLens20M [^1] https://grouplens.org/datasets/movielens/ This is a widely used benchmark dataset Harper & Konstan (2015) for evaluating personalization and recommendation algorithms. Interaction features are movie name, genre, and rating.
  • Google Local Review Dataset [^2] https://datarepo.eng.ucsd.edu/mcauley_group/gdrive/googlelocal/ Li et al. (2022); Yan et al. (2023) This dataset contains review information on Google Maps up to Sep 2021. We used New York’s data in our experiments. Interaction features are place name, place category, rating, and review.
  • Amazon Review Dataset [^3] https://nijianmo.github.io/amazon/index.html He & McAuley (2016) This is a series of product review datasets crawled from Amazon.com. We used the ”Movie_and_TV” category for evaluation. Interaction features are product title, product category, rating, and review summary, which is a short opinion summary written by the user.
We generated training and test examples by applying a sliding window over each user’s feature sequences (sorted by timestamps). This creates inputs containing N items, with the subsequent item as the label. To ensure model evaluation consistency, the final example (representing the most recent interaction) is used as the test example for each user and is never included in the training set.
To address the diverse evaluation needs, we accessed model performance on three distinct types of tasks:
@@
  • Favorite genre or category prediction The model predicts a user’s favorite genre or category based on a sequence of items (e.g., predicting favorite movie genre given a sequence of movie IDs). The favorite is the one with the highest frequency in the user’s history. This simple task serves as a proof-of-concept, demonstrating User-LLM’s ability to extract and understand user preferences from historical interactions.
  • Multimodal review generation This task extends beyond simple predictions. Here, the model must generate actual reviews, utilizing multimodal input features (e.g., movie ID, genre, and rating). These features are encoded to create embeddings that guide the generation of a user’s potential review for a given item.
@@ 4.2 Baselines and Experiment Setup
To assess the effectiveness of User-LLM, we compare it against the following baselines:
  • Dual Encoder(DualEnc) An embeddings-based two-tower model that is widely used for large-scale recommendation systems and natural language tasks.
  • Bert4Rec The state-of-the-art language model powered recommendation model BERT4Rec Sun et al. (2019). To ensure fair comparison, Bert4Rec employs similar transformer parameters as our user encoder.
  • TextPrompt(TP) An LLM contextualized with user history via text-prompt-based finetuning on the raw text user history.
@@
Our experiments primarily used an Autoregressive Transformer for embedding generation, with ablation studies comparing performance against a Dual Encoder architecture. For co-training, we explored four different training strategies described in Section 3.2.3 and evaluated the impact of using a pretrained encoder versus a randomly initialized one.
For all text-prompt-based LLM finetuning and User-LLM experiments, we kept the total number of examples being processed by the model the same for comparison. For Dual Encoder and Bert4Rec baselines, we did hyper-parameter tuning and report the best performance results.
@@
Our experiments utilized PaLM-2 XXS Anil et al. (2023) as the pretrained LLM. The user encoder is a $6$-layer $128$-dimensional transformer with $8$ attention heads. For Perceivers, we constructed the module with $6$ cross-attention layers and a query dimension of $16$. We used $rank=8$ for all LoRA tuning experiments.
4.3 Overall Performance Comparison
@@
This subsection compares model accuracy across various tasks and baselines. All reported User-LLM experiments utilize pretrained encoders and finetune all three components (encoder, projection layers, and LLM) during cotraining, unless stated otherwise.
4.3.1 Next Item Prediction
Table 2 shows that User-LLM outperforms the two non-LLM baselines on MovieLens and Google Local review datasets for next item prediction. However, Bert4Rec excels on the sparse Amazon Review dataset, likely due to its bidirectional architecture, which enables more robust representation learning from limited training examples. In contrast, our autoregressive (unidirectional) approach relies on larger datasets for effective user modeling. This suggests that Amazon Review dataset may not provide sufficient information for the autoregressive approach to learn effectively for next item prediction tasks.
Dataset Rec Baseline User-LLM
DualEnc Bert4Rec
MovieLens20M @1 0.044 0.038 0.054
@5 0.135 0.135 0.164
@10 0.206 0.158 0.243
Google Review @1 0.005 0.005 0.015
@5 0.012 0.019 0.052
@10 0.023 0.033 0.071
Amazon Review @1 0.021 0.034 0.037
@5 0.026 0.051 0.047
@10 0.031 0.062 0.051
Table 2: User-LLM v.s. DualEnc & Bert4Rec baselines for next item prediction. Note that we use full finetune for User-LLM.
Fig. 3 compares User-LLM and text-prompt-based LLM finetuning on MovieLens20M (more results in Appendix).
Varying Context Length When trained on sequences with predefined lengths (e.g., 50, 100, or 200 items) and all parameters tunable (solid lines in Fig. 3), User-LLM shows slightly lower performance than text-prompt-based finetuning with shorter sequences (50 items). However, User-LLM significantly outperforms the text-prompt-based approach when sequences become longer. This is likely because next-item prediction heavily relies on the LLM’s ability to memorize specific item names (e.g., movie titles). Text prompt excels in this type of task by providing the LLM with numerous names or titles to memorize. As sequences become longer, they introduce diverse and potentially noisy information, making pure memorization less effective. It is a known limitation that LLMs are less effective with long input context Liu et al. (2023c). In contrast, User-LLM’s ability to effectively distill relevant user information from these complex long sequences enhances LLM’s understanding and leads to better performance.
With increasing input sequence length, both User-LLM and text-prompt-based approach experience performance degradation. This is likely due to the significantly reduced number of training examples available (13.8M, 10.6M, and 7.0M for lengths 50, 100, and 200, respectively). While the autoregressive encoder impacts model performance due to insufficient training data, User-LLM’s performance decline is notably less severe compared to the text-prompt-based approach.
Computation Efficiency Notably, the computation cost and memory requirements of text-prompt-based approaches become increasingly prohibitive with increasing input sequence length. Table 3 highlights User-LLM’s computational efficiency. The user encoder distills ID-based user activity sequences into compact user embeddings, enabling the LLM to process a fixed-length $32$-token input query regardless of input sequence length. Conversely, the text-prompt-based approach requires LLM input token lengths to be scaled proportionally with the input sequence.
Specifically, let the number of LLM parameters be $N$, encoder parameters be $N_{e}$, and Perceiver parameters be $N_{p}$. As $N_{e}\ll N$ and $N_{p}\ll N$, we can ignore computational costs associated with the encoder and Perceiver. Assuming batch size $B$, training steps $S$, and input length (in tokens) $L$, the total number of training tokens is $D=BSL$. Applying heuristic $FLOPs\approx 6ND$ Kaplan et al. (2020), User-LLM shows up to 78.1X FLOPs reduction when the sequence length reaches 200. This demonstrates that User-LLM provides a more computationally efficient and practically feasible solution for long sequence modeling.
Figure 3: User-LLM v.s. TextPrompt(TP) baseline for next item prediction on MovieLens20M (varying sequence lengths, unfrozen or frozen LLM). Unfrozen LLM: full finetune for TP model, training strategy Full for User-LLM. Frozen LLM: LoRA parameters tuning for TP model, training strategy Enc for User-LLM.
@@ 4.3.1 Next Item Prediction
Table 3: LLM input token counts for User-LLM v.s. TextPrompt(TP) and FLOPs reduction achieved by User-LLM. FLOPs reduction refers to FLOPs(TP) / FLOPs(User-LLM).
Frozen LLM User-LLM excels in contextualizing LLMs while preserving their original knowledge. As described in Section 3.2.3, User-LLM allows the LLM to remain frozen (e.g., training strategy Enc), preventing catastrophic forgetting issues. Fig. 3 (dotted lines) demonstrates that User-LLM with a frozen LLM surpasses the accuracy of text-prompt-based LoRA tuning. This observation suggests that performance gains in text-prompt-based full finetuning may largely result from overfitting, leading to significant knowledge loss within the LLM.
Furthermore, User-LLM with a frozen LLM also outperforms models with unfrozen LLMs (both User-LLM and text-prompt-based approach). This highlights the value of the foundation knowledge encoded in LLMs for improving downstream task performance. By utilizing user embeddings, User-LLM enables efficient and stable LLM contextualization, effectively mitigating the drawbacks of text-prompt-based methods.
4.3.2 Deep User Understanding
Even though user encoders are pretrained on next item prediction tasks, the extracted user embeddings encapsulate generalized user context. Leveraging these embeddings, User-LLM demonstrates great generalization capability to diverse personalization tasks, as shown in Table 4. These tasks require deep user understanding, demonstrating User-LLM’s effectiveness in understanding user intent and preferences from interaction data.
User-LLM outperforms the Dual Encoder baseline, which is trained to predict genre/category ID based on a sequence of user history item IDs. Despite its encoder’s pretraining on next-item prediction, User-LLM surpasses the specialized Dual Encoder across all three datasets. Compared to the TextPrompt baseline, User-LLM excels on Amazon Review and achieves similar performance on MovieLens and Google Local Review, with a significantly smaller context window (16 times smaller), emphasizing its efficiency.
User-LLM also demonstrates its strength in the multimodal task of review generation based on the item name, category, and user rating. It exceeds the TextPrompt baseline on both Amazon Review and Google Local Review datasets. We attribute this success to User-LLM’s ability to leverage dense representations of user history, an advantage particularly evident in tasks with long inputs where text-prompt-based LLM finetuning often struggles. Note that Dual Encoder cannot handle such text generation tasks.
Dataset Task Baseline User-LLM
DualEnc TP
MovieLens 20M FavGenre 0.779 0.788 0.785
Google Review FavCat 0.812 0.887 0.862
ReviewG - 10.20 11.72
Amazon Review FavCat 0.855 0.885 0.890
ReviewG - 22.82 26.38
Table 4: Favorite genre/category generation and review generation. Metrics: Recall@1 for favorite genre/category prediction tasks and ROUGE-Lsum for review generation.
4.4 Efficiency
Parameter efficiency Embedding-LLM requires fewer tunable parameters to achieve competitive performance. Table 5 shows the model accuracy across multiple training strategies. Notably, the Enc approach (tuning encoder and projection layers and freezing the LLM) achieved comparable task accuracy to full finetuning while requiring significantly fewer tuned parameters. This is because it effectively leverages the pretrained weights of both the user encoder and LLM, minimizing the need for extensive parameter updates during training. Thus, our approach offers a parameter-efficient solution for personalizing LLMs, reducing computational demands and facilitating model tuning on resource-constrained devices.
Dataset Task Full LoRA Enc Proj
MovieLens 20M NextItem 0.054 0.053 0.052 0.040
FavGenre 0.785 0.784 0.783 0.750
Google Review NextItem 0.015 0.016 0.015 0.015
FavCat 0.862 0.844 0.844 0.615
ReviewG 11.72 11.64 11.76 9.61
Amazon Review NextItem 0.037 0.028 0.028 0.014
FavCat 0.890 0.887 0.885 0.850
ReviewG 26.38 24.56 24.30 19.04
Table 5: Comparison between different training strategies. Metric: Recall@1.
Inference efficiency Furthermore, User-LLM is inference efficient. Compared to text-prompt-based approaches that require multiple tokens to represent each event in the user history, our embedding-based approach effectively condenses event information into dense representations, requiring only a single token per event. The integration of Perceiver further reduces the total number of tokens for long user histories, resulting in even faster inference speeds and making our approach ideal for real-time applications. Table 6 shows that User-LLM and its Perceiver variant maintain competitive performance while requiring a smaller number of tokens and, as a result, significantly reduce the computation complexity in the attention modules in LLM.
Task User-LLM User-LLM +Perceiver

### v2

@@ 3.3 Training Framework
User-LLM offers flexible training strategies for Encoder-LLM finetuning to tailor its performance for diverse use cases. To thoroughly explore User-LLM’s capabilities, we investigated four strategies targeting different model components:
  • $\circ$
@@
  • $\circ$
    LoRA: Finetuning the LLM with LoRA (Hu et al. 2022), along with the user encoder and projection layers, offers parameter efficiency preventing catastrophic forgetting and maintaining the LLM’s core knowledge.
  • $\circ$
@@
3.4 Efficiency
Compared to text-prompt-based methods, User-LLM offers substantial efficiency gains. First, it leverages pretrained weights of both the user encoder and LLM, improving model convergence speed with less trainable parameters. Additionally, User-LLM condenses user activities into dense representations, requiring only a single token per event. This frees up the LLM’s context window, leading to improved inference efficiency. Furthermore, User-LLM leverages cross-attention, maintaining user embeddings at a smaller dimension than the LLM’s model dimension, which significantly reduces computation during both training and inference.
To further optimize inference efficiency, User-LLM integrates Perceiver (Jaegle et al. 2021) units into its projection layers. Perceiver is a Transformer-based architecture that utilizes a trainable latent query to extract relevant information from input embeddings through cross-attention. In User-LLM, Perceiver effectively compresses the user embeddings into a compact format using a latent query with shorter sequence length. This compression reduces the number of tokens needed to represent the user history, further freeing up the LLM’s context window. This, along with Perceiver’s ability to distill insights from noisy contexts, makes User-LLM ideal for practical applications with extensive user histories.
@@ 4.1 Datasets and Tasks
We evaluated our approach on three widely recognized datasets: MovieLens20M (Harper and Konstan 2015), Amazon Review (He and McAuley 2016), and Google Review (Li, Shang, and McAuley 2022; Yan et al. 2023). Notably, Amazon Review dataset is well known for its high sparsity and variability and contains a significantly smaller number of training and evaluation examples.
We generated training and test examples by applying a sliding window over each user’s feature sequences (sorted by timestamps). This creates inputs containing N items, with the subsequent item as the label. To ensure model evaluation consistency, the final example (representing the most recent interaction) is used as the test example for each user and is never included in the training set. Table 1 summarizes datasets statistics after processing.
To address the diverse evaluation needs, we accessed model performance on three distinct types of open-text-generation tasks. These tasks serve as a proof-of-concept, demonstrating User-LLM’s ability to understand user preferences from historical interactions and provide peronslized response accordingly.
  • $\circ$
@@
  • $\circ$
    Favorite genre or category prediction The model predicts a user’s favorite genre or category based on a sequence of items (e.g., predicting favorite movie genre given a sequence of movie IDs). The favorite is the one with the highest frequency in the user’s history.
  • $\circ$
    Review generation This task extends beyond simple predictions. Here, the model must generate actual reviews, utilizing multiple input features (e.g., movie ID, genre, and rating). These features are encoded to generate embeddings that guide the generation of a user’s potential review for a given item.
Datasets #Users #Items #Train Seq Len
MovieLens20M 82,977 27,280 13,821,405 50
Google Review 80,080 270,721 3,387,375 50
Amazon Review 5,756 177,978 357,258 50
Table 1: Dataset statistics. #Test is equal to #Users.
@@ 4.2 Baselines and Experiment Setup
Our experiments used an Autoregressive Transformer for embedding generation. For co-training, we explored four different training strategies described in Section 3.3 and evaluated the impact of using a pretrained encoder versus a randomly initialized one.
@@
We utilized PaLM-2 XXS (Anil et al. 2023) as the pretrained LLM. The user encoder is a $6$-layer $128$-dimensional transformer with $8$ heads. We constructed the Perceivers module with $6$ cross-attention layers and a query dimension of $16$. We used $rank=8$ for all LoRA tuning experiments.
To assess the effectiveness of User-LLM, we compare it against the following two baselines:
  • $\circ$
    Text Prompt (TP) contextualize LLM by finetuning it with raw text input derived from a user’s timeline.
  • $\circ$
    Text Summarization (TS) contextualize LLM by finetuning it with text Summarization derived from a user’s timeline. In particular, we first use a Gemini-M model to generate summarizations for raw text inputs of user history, and then finetune the PaLM-2 XXS towards different tasks. This is a time consuming process, and due to resource limitation, we only conduct comparison on Amazon Review datasets.
4.3 Performance and Efficiency Comparison
Dataset Task Metrics Text-Prompt User-LLM Relative Improvement FLOPs Reduction
MovieLens 20M NextItem Recall@5 0.140 0.154 10.00% 21.9X
FavGenre Accuracy 0.787 0.787 0.00% 21.9X
Google Review NextItem Recall@5 0.057 0.052 -8.77% 12X
FavCategory Accuracy 0.741 0.862 16.33% 12X
ReviewGen Rouge 10.20 11.72 14.90% 16X
Amazon Review NextItem Recall@5 0.042 0.047 11.90% 16X
FavCategory Accuracy 0.885 0.890 0.57% 16X
ReviewGen Rouge 22.82 26.38 15.59% 16X
Table 2: User-LLM vs. Text-Promp for a variety of tasks on efficiency and performance. Note that we use full finetune for User-LLM. Relative improvement is calculated as (User-LLM - TP) / TP * 100%.
@@ 4.3 Performance and Efficiency Comparison
This subsection compares model accuracy and efficiency across various tasks and baselines. All reported User-LLM experiments utilize pretrained encoders and finetune all three components (encoder, projection layers, and LLM) during cotraining, unless stated otherwise.
Overall Performance
User-LLM demonstrates significant efficiency gains (Table 2), achieving a remarkable 12-21.9X reduction in FLOPs compared to the Text Prompt baseline. Moreover, User-LLM consistently outperforms the Text Prompt baseline on most tasks across all three datasets, with performance improvements up to 16.33%. The exception is the Google review next item task, where User-LLM exhibits slightly lower performance.
While pretrained on next item prediction tasks, the user encoders generates embeddings encapsulating generalized user context, enabling User-LLM to generalize well to diverse personalization tasks (e.g., favorite genre/categroy prediction, review generation) (Table 2). These tasks require deep user understanding, demonstrating User-LLM’s effectiveness in understanding user preferences from interaction data.
Figure 3: User-LLM v.s. Text Prompt (TP) v.s. Text Summarization (TS) for 3 tasks on Amazon Review dataset. User-LLM achieves better performance with much fewer tokens.
Fig. 3 compares the performance of User-LLM against Text Prompt (TP) and Text Summarization (TS) baselines on three tasks using the Amazon Review dataset: Next Item Prediction, Favorite Category Prediction, and Next Review Generation. The x-axis represents the number of tokens used, highlighting the efficiency of each method. Notably, User-LLM consistently achieves superior performance while utilizing significantly fewer tokens. This efficiency is evident in all three tasks, indicating the effectiveness of directly embedding user history as a distinct modality. User-LLM requires fewer tokens to achieve equal or better performance, suggesting its ability to extract and leverage more relevant information from user data.
Varying Context Length
To understand the impact of input length on performance and efficiency, we trained User-LLM with varying predefined sequence lengths (50, 100, and 200 items) on MovieLens20M. Both frozen and tunable LLMs were evaluated and compared against a text-prompt baseline fine-tuned under the same computational constraints.
Table 4 highlights the significant computational advantage of User-LLM. As the input sequence length increases, the required number of input tokens for the text-prompt approach scales proportionally (700, 1350, and 2500 tokens for lengths 50, 100, and 200 respectively). This leads to a substantial increase in computational cost and memory demands. In contrast, User-LLM leverages a user encoder that compresses user activity sequences into compact embeddings. This allows the LLM to operate on a fixed-length input of only 32 tokens, regardless of the input sequence length. To illustrate the efficiency gain, we can approximate FLOPs using the heuristic $FLOPs\approx 6ND$ (Kaplan et al. 2020), where $N$ represents the number of LLM parameters, and $D$ represents the total number of training tokens ($D=BSL$, with batch size $B$, training steps $S$, and input length $L$). Given that the encoder and Perceiver parameters ($N_{e}$ and $N_{p}$ respectively) are significantly smaller than the LLM parameters ($N_{e}\ll N$ and $N_{p}\ll N$), their computational costs are considered negligible. Using this approximation, User-LLM achieves a remarkable FLOPs reduction compared to the text-prompt baseline, reaching up to 78.1X when the input sequence length reaches 200.
Frozen LLM
Furthermore, User-LLM excels at integrating LLMs while preserving their inherent knowledge. By utilizing a training strategy (Enc) that keeps the LLM frozen, we prevent catastrophic forgetting and maintain the LLM’s pretrained capabilities. As demonstrated in Table 4, User-LLM with a frozen LLM surpasses the performance of text-prompt-based methods utilizing both full fine-tuning and LoRA tuning. Impressively, it achieves comparable performance to User-LLM trained with an unfrozen LLM. These findings highlight User-LLM’s ability to effectively contextualize user preferences while safeguarding the LLM’s pretrained knowledge, effectively mitigating the drawbacks associated with text-prompt-based methods.
Parameter efficiency
User-LLM requires fewer tunable parameters to achieve competitive performance. Table 3 shows the model accuracy across multiple training strategies. Notably, the Enc approach (tuning encoder and projection layers and freezing the LLM) achieved comparable task accuracy to full finetuning while requiring significantly fewer tuned parameters. This is because it effectively leverages the pretrained weights of both the user encoder and LLM, minimizing the need for extensive parameter updates during training. Thus, our approach offers a parameter-efficient solution for personalizing LLMs, facilitating model tuning on resource-constrained devices.
Dataset Task Metric Full LoRA Enc Proj
MovieLens 20M NextItem Recall@1 0.055 0.051 0.050 0.023
FavGenre Recall@1 0.787 0.787 0.786 0.743
Google Review NextItem Recall@1 0.015 0.016 0.015 0.015
FavCat Recall@1 0.862 0.844 0.844 0.615
ReviewG Rouge 11.72 11.64 11.76 9.61
Amazon Review NextItem Recall@1 0.037 0.028 0.028 0.014
FavCat Recall@1 0.890 0.887 0.885 0.850
ReviewG Rouge 26.38 24.56 24.30 19.04
Table 3: Comparison between different training strategies.
@@
Performance improvement Unfrozen LLM 1.10X 1.23X 1.49X
Frozen LLM 2.78X 2.97X 3.27X
Table 4: LLM input token counts for User-LLM v.s. TextPrompt(TP). FLOPs reduction and performance improvement achieved by User-LLM. FLOPs reduction = FLOPs(TP) / FLOPs(User-LLM). Performance improvement = Recall@1(User-LLM) / Recall@1(TP).
Task User-LLM +Perceiver
@@ 4.4 Ablation Study
Figure 4: Converged gating parameters (Alpha) of different cross-attention layers indicate upper LLM transformer layers attend more signifcantly to encoder output than lower layers in next item prediction models on all three datasets.
Benefit of pretraining
We study the benefits of pretraining user encoders for downstream tasks. Results presented in columns 3 to 4 in Table 6 demonstrate that models utilizing pretrained encoders consistently outperform those with randomly initialized encoders across all tasks, suggesting that pretraining enables encoders to capture user preference effectively, providing valuable context to LLMs during cotraining.
Soft Prompt v.s. Cross Attention
We compare the performance of User-LLM with an alternative encoder-LLM integration strategy, soft-prompt (Lester, Al-Rfou, and Constant 2021). We set the task prompt’s length to $10$ and utilized a pretrained Autoregressive encoder in our experiment. As shown in the last column of Table 6, the cross-attention-based approach generally outperforms the soft-prompt approach. This advantage is particularly significant in the two review generation tasks, suggesting that the cross-attention mechanism enables LLMs to better extract and utilize the rich user information encoded within user embeddings. This highlights the value of cross-attention for tasks demanding a nuanced understanding of human intent.
Dataset Task cross-attention soft prompt
PT RD
MovieLens 20M NextItem 0.055 0.043 0.047
FavGenre 0.787 0.782 0.787
Google Review NextItem 0.015 0.015 0.013
FavCat 0.862 0.853 0.822
ReviewG 11.72 11.43 9.06
Amazon Review NextItem 0.037 0.026 0.031
FavCat 0.890 0.882 0.894
ReviewG 26.38 23.12 25.80
Table 6: Experiments on different model fusing architectures, encoders, and the benefit of pretraining. Report Recall@1 for next item prediction and favorite genre/category prediction, ROUGE-Lsum for review generation. Full finetuning is used throughout the experiments. PT: Pretrained encoder. RD: Randomly initialized encoder. soft-prompt: Connect user embeddings with LLM via soft-prompt.
Gated Cross-attention
We further assess cross-attention in User-LLM by employing gated cross-attention layers that include a trainable gating parameter $\alpha$ governing the amount of cross-attention to be added to the language model stack. Fig. 4 shows the converged values of gating parameters of 18 gated cross-attention layers for next item prediction tasks on MovieLens20M, Amazon Review and Google Local Review. The $\alpha$ values suggest that the upper LM layers attend more significantly to encoder output compared to the lower layers. Since the encoder output is user embedding that typically captures high-level semantics such as user attributes and interests, our result suggests that the upper layers in User-LLM are likely to process such semantic information, hence attending more to encoder output. Interestingly, previous work by Geva and others (Geva et al. 2021) showed that lower layer keys in transformer correlate more with shallow surface features, whereas upper layer keys capture semantic topics. These analyses provide a plausible explanation why cross-attention outperforms soft-prompt in User-LLM.
5 Conclusion and Future Work
In this paper, we introduced User-LLM, a framework for contextualizing LLMs through user embeddings. These embeddings, derived from self-supervised pretraining on diverse user interactions, capture hidden user preferences and their evolution. By integrating these embeddings with LLMs through cross-attention, User-LLM empowers LLMs to adjust dynamically to user contexts.
Our comprehensive evaluation on MovieLens, Amazon Review, and Google Local Review datasets demonstrated that User-LLM consistently outperforms text-prompt-based approaches on tasks requiring deep user understanding, particularly those involving long user interaction histories. User-LLM’s computational efficiency and ability to preserve LLM knowledge further make it particularly well-suited for practical applications where real-time personalization is crucial.
Future research directions include optimizing user embedding generation through advanced pretraining techniques and investigating the alignment between user embeddings and the language model for even deeper user context understanding. Additionally, training User-LLM on a diverse range of tasks could further enhance its generalization abilities and adaptability across a broader spectrum of user scenarios. By addressing these aspects, User-LLM has the potential to unlock the full power of LLMs for building truly personalized and adaptive AI agents.
Appendix A More details about Datasets
  • MovieLens20M [^1] https://grouplens.org/datasets/movielens/ A widely used benchmark dataset (Harper and Konstan 2015) for evaluating personalization and recommendation algorithms. Interaction modalities are movie name, genre, and rating.
  • Google Local Review Dataset [^2] https://datarepo.eng.ucsd.edu/mcauley_group/gdrive/googlelocal/ (Li, Shang, and McAuley 2022; Yan et al. 2023) This dataset contains review information on Google Maps up to Sep 2021. We used New York’s data in our experiments. Interaction modalities are place name, place category, rating, and review.
  • Amazon Review Dataset [^3] https://nijianmo.github.io/amazon/index.html (He and McAuley 2016) This is a series of product review datasets crawled from Amazon.com. We used the ”Movie_and_TV” category for evaluation. Interaction modalities are product title, product category, rating, and review summary (a short opinion summary written by the user).
Appendix B Training Hyperparameters
Dataset Task Learning rate Weight decay Batch size Training steps
MovieLens 20M EncoderPretrain 1e-1 1e-1 32,768 10,000
NextItem 1e-3 1e-3 8,192 10,000
FavGenre 1e-3 1e-3 1,024 5,000
Google Review EncoderPretrain 1e-4 1e-1 16,384 10,000
NextItem 1e-3 1e-1 2,048 5,000
FavCat 1e-3 1e-1 2,048 5,000
ReviewG 1e-3 1e-1 1,024 5,000
Amazon Review EncoderPretrain 1e-4 1e-1 12,800 7,000
NextItem 1e-3 1e-1 8,192 10,000
FavCat 1e-3 1e-1 1,024 5,000
ReviewG 1e-3 1e-1 2,048 5,000
Table 7: Training hyperparameters.
Appendix C Encoder Architectures
User-LLM uses a Transformer-based user encoder to generate user embeddings from ID-based feature sequences. To effectively fuse information from user interaction data, we explore and compare two fusion architectures: early fusion and late fusion. Following the example shown in Fig. 5, each item in the user interaction timeline has three features: name, rating, and category. Each feature has its own vocabulary. Items from each feature are mapped to an integer ID and have their own embedding representations. In an early fusion architecture, the $i_{th}$ item’s three features (name, rating, and category) are combined into a fused embedding $f_{i}$. The sequence of fused embeddings is then fed into the user encoder to generate user embeddings. On the other hand, a late fusion architecture employs three separate feature encoders to process each feature independently. The resulting embeddings are combined to a single user embedding at a later stage using fusion layers.
Autoregressive Transformer As described in Section 3.2, User-LLM employs an Autoregressive Transformer as the early fusion encoder. Illustrated in Fig. 2, this model receives a sequence of user activities, where each activity is represented by multiple features (name, rating, and category). Individual features are first embedded, then concatenated and processed by a Transformer decoder. The Autoregressive Transformer Encoder generates a user embedding for each input item in the user’s timeline.
Compared to Autoregressive Transformers, Dual Encoders offer several advantages. The separate feature encoders within user towers allow independent processing of features, handling features with varying sequence lengths or misaligned timestamps and potentially enhancing information extraction. The configurable fusion layers provide flexibility in the number of generated embeddings. Additionally, Dual Encoders support diverse pretraining tasks beyond next item prediction, as user and label towers can process different features (e.g., name for the user tower and category for the label tower). However, Dual Encoders may be less effective at modeling the complex sequential patterns often present in user timelines. The optimal encoder choice depends on dataset characteristics, the desired level of contextual understanding, and a balance between computational cost and potential performance gain.
Table 9 compares the performance of using Autoregressive Transformer (AR) and Dual Encoder (DE) on three datasets and eight tasks, without computation constraint. AR consistently outperforms DE, across various training strategies and using both pretrained or randomly initialized encoders. Table 10 further analyzes the two encoders’ performance when incorporating short-term history into the LLM prompt.
Appendix D Integration Approaches
As mentioned in 4.3, we explored the integration of user embeddings and LLMs in User-LLM through two approaches: cross-attention and soft-prompt. In the cross-attention method, the output embeddings from the pretrained user encoder are cross-attended with the LLM’s intermediate text representations, similar to how Flamingo (Alayrac et al. 2022) works. On the other hand, the soft-prompt approach leverages user embeddings from the pretrained encoder as a soft prompt for the LLM. This soft prompt is prepended to the regular text prompt embedding, providing additional context about the user’s interests and preferences.
Table 11 presents a detailed comparison between cross-attention and soft-prompt approaches with various training strategies across all datasets and tasks.